OpenStudy (anonymous):

the popuation density of New Mexico increased at a relatively constant rate from 1980 to 1999. in 1985, the population density was about 11.62 people per square mile. in 1999, the population density was about 14.25 people per square mile. Write an equation that gives the population density (in square mile) as a function of the number of years since 1980. what was the population density in 1990.

OpenStudy (anonymous):

slope intercept form

OpenStudy (anonymous):

slope intercept form: y= mx+ b x, in years since 1980 (5, 11.62) (19, 14.25) First, find the slope (m), using the formula \[y1-y2 \over x1 - x2\] then plug in x and y from either one of the coordinates given into the y=mx+b formula. When you have found that, you want to find the pop. density for 1990, where x is 10. Plug it into your system and you should be able to find the pop.density for that year easy peasy.

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

would the slope be \[\frac{ 263 }{1400}\] then

OpenStudy (anonymous):

that's exactly right

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so it would be \[14.25=\frac{ 263 }{ 1400 }(19)+m\]